a television addict can buy either two televisions and three video recorders for $1750 or four televisions and one video recorder for $1250. find the cost of one of each.
Answer by checkley71(8403)

2T+3V=1750 OR 2T=1750-3V
4T+V=1250 REPLACE 4T BY 2(1750-3V)=3500-6V
3500-6V+V=1250
-5V=1250-3500
-5V=-2250
V=-2250/5
V=$450 FOR THE VIDEO.
2T+3(450)=1750
2T+1350=1750
2T=1750-1350
2T=400
T=400/2
T=200 FOR THE TV.
PROOF
4*200+450=1250
800+450=1250
1250=1250